OpenStudy (patrickweewoos):

Which equation models this problem? Chelsea paid $54.90 for nine bath towels. What is the unit price, p, per towel?

OpenStudy (patrickweewoos):

@perl and @iGreen

OpenStudy (anonymous):

divide again..

OpenStudy (igreen):

You can use the equation: \(\dfrac{54.90}{9} = p\) or \(9p = 54.90\). @PatrickWeeWoos

OpenStudy (anonymous):

54.90 / 9 = ?

OpenStudy (patrickweewoos):

A.)p/9=54.90 B.)54.90p=9 C.)9p=54.90 D.)p/54.90=9

OpenStudy (patrickweewoos):

@iGreen what do you think

OpenStudy (igreen):

I already listed it..lol.

OpenStudy (patrickweewoos):

ABC or D

OpenStudy (igreen):

I said: You can use the equation: 54.90/9=p or 9p=54.90.

OpenStudy (patrickweewoos):

so its C

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes its C

OpenStudy (igreen):

Yep. @PatrickWeeWoos